I am having a REALLY hard time with this:
Complete the given table of values and graph the equation.
y=x^2-5x+4
The given table has nothing under y.
Under x looking downward is 0,1,2,3,4.
Your help would be greatly appreciated!
Thanks~
ac

Answer by Edwin McCravy(20086)

I am having a REALLY hard time with this:
Complete the given table of values and graph the equation.
y=x^2-5x+4
The given table has nothing under y.
Under x looking downward is 0,1,2,3,4.
Your help would be greatly appreciated!
Thanks~
ac

 x   y
 0   ?
 1   ? 
 2   ?
 3   ?  
 4   ?

First substitute x = 0
y = x² - 5x + 4
y = 0² - 5(0) + 4
y = 0 - 0 + 4
y = 4

so we write 4 in the y column next to 0
in the x column, and we have

 x   y
 0   4
 1   ?
 2   ?
 3   ? 
 4   ?

Next substitute x = 1
y = x² - 5x + 4
y = 1² - 5(1) + 4
y = 1 - 5 + 4
y = 0

so we write 0 in the y column next to 1
in the x column, and we have

 x   y
 0   4
 1   0
 2   ? 
 3   ?
 4   ?   

Next substitute x = 2
y = x² - 5x + 4
y = 2² - 5(2) + 4
y = 4 - 10 + 4
y = -2

so we write -2 in the y column next to 2
in the x column, and we have

 x   y
 0   4
 1   0
 2  -2
 3   ?
 4   ?   

Next substitute x = 3
y = x² - 5x + 4
y = 3² - 5(3) + 4
y = 9 - 15 + 4
y = -2

so we write -2 in the y column next to 3
in the x column, and we have

 x   y
 0   4
 1   0
 2  -2
 3  -2
 4   ?

Next substitute x = 4
y = x² - 5x + 4
y = 4² - 5(4) + 4
y = 16 - 20 + 4
y = 0

so we write 0 in the y column next to 4
in the x column, and we have

 x   y
 0   4
 1   0
 2  -2
 3  -2
 4   0

Now we plot the four points
(0,4), (1,0), (2,-2), (3,-2), and (4,0)



Then we draw a smooth curve through those 
five points:



Edwin